OpenStudy (nptweezy):

Solve the equation for x. 50 = –6x + 4x – 8 (UPDATED)

OpenStudy (anonymous):

Add 10 to both sides. Combine terms. Divide by sides by the number attached to the variable.

OpenStudy (anonymous):

*both sides

OpenStudy (nptweezy):

its 0 @Nanoman

OpenStudy (uri):

3x+6x=-10+10 9x=0 x=0/9 x=0 No solution.

OpenStudy (nptweezy):

@uri 0 was the solution

OpenStudy (uri):

Lol yea :p x=0.

OpenStudy (nptweezy):

@Nanoman how do I do this one

OpenStudy (uri):

Wait one Question at a time,@NPTWEEZY

OpenStudy (anonymous):

Yes, the answer is 0.

OpenStudy (anonymous):

You have another problem?

OpenStudy (uri):

@Nanoman He updated it..But only on question at a time is allowed so close it if you don't want to get suspended @NPTWEEZY